# SubIQ: Inverse Soft-Q Learning for Offline Imitation with Suboptimal Demonstrations

### v1

@@ 3.2 Regularization with Reward Reference
where $\alpha>0$ is a weight parameter for the reward regularizer term. With (5), the goal is to find a policy functions that match with the occupancy measure of the expert policies. Simultaneously, it ensures that the learning rewards are close to the pre-assigned rewards, aiming to guide the learning policy towards replicating expert demonstrations.
Even though (5) can be directly solved to recover rewards, prior research suggests that transforming (5) into the Q-space will enhance efficiency. We delve into this approach in the next section.
@@
3.3 Tractable Lower-bound on Inverse Soft-Q
As discussed earlier, there is a one-to-one mapping between any reward function $r$ and a function $Q$ in the Q-space. Thus, the maximin problem in (5) can be equivalently transformed into the $Q$-space as:
$\max_{Q}$ $\min_{\pi}\Big{\{}{\mathcal{H}}(Q,\pi)\stackrel{{\scriptstyle def}}{{=}}\sum_{i\in[N]}w_{i}\mathbb{E}_{\rho^{i}}[{\mathcal{T}}^{\pi}[Q](s,a))]$
$-\mathbb{E}_{\rho_{\pi}}[{\mathcal{T}}^{\pi}[Q](s,a))]+{\mathbb{E}_{\rho_{\pi}}[\log\pi(s,a)]}$
@@ 3.3 Tractable Lower-bound on Inverse Soft-Q
${\mathcal{T}}^{\pi}[Q](s,a)=Q(s,a)-\gamma\mathbb{E}_{s^{\prime}}[V^{\pi}(s^{\prime})]$
$V^{\pi}(s)=\mathbb{E}_{a\sim\pi(a|s)}[Q(s,a)-\log\pi(a|s)]$
In the context of single-expert-level, (Garg et al., 2021) demonstrated that the objective function in the $Q$-space is concave in $Q$ and convex in $\pi$, implying that the maximin problem always has a unique saddle point solution. Moreover, the inner minimization problem over $\pi$ will yield a unique closed-form solution, enabling the transformation of the maximin problem into a non-adversarial concave problem within the Q-space. This makes the inverse soft-Q problem advantageous to handle. However, in our context, these results do not hold. Let us first state the following proposition, which asserts that ${\mathcal{H}}(Q,\pi)$ is not necessarily concave in $\pi$.
@@
${\mathcal{H}}(Q,\pi)$ is concave in $Q$ but is not convex in $\pi$.
In general, we can see that the first and second term of (7) are convex in $Q$, but the reward regularizer term, which can be written as $\alpha\mathbb{E}_{\rho^{u}}\Big{[}Q(s,a)-\overline{r}(s,a)-\mathbb{E}_{s^{\prime}\sim P(.|s,a)}\mathbb{E}_{a^{\prime}\sim\pi(.|s^{\prime})}(Q(s^{\prime},a^{\prime})-\log\pi(s^{\prime},a^{\prime})))^{2}\Big{]}$, is not concave (or concave) in $\pi$ (details are shown in the Appendix). The property indicated in Proposition 3.1 implies that the maximin problem within the $Q$-space $\max_{Q}\min_{\pi}J(Q,\pi)$ may not have a unique saddle point solution and would be more challenging to solve, compared to the original inverse IQ-learn problem.Additionally, to circumvent non-adversarial training, (Garg et al., 2021) demonstrated that the inner minimization problem in $\pi$ always has a closed-form solution such that $\pi^{Q}=\text{argmax}_{\pi}~{}~{}V^{\pi}(s)$ for all $s\in S$. This result is essential to convert the minimax problem into a concave maximization problem in $Q$. Unfortunately, this result does not hold with the new objective function in (7), as shown below:
@@
[…]
To overcome the above challenges, our approach involves constructing a more tractable objective function that lower-boundedly approximates the objective of (7). To this end, let us first define $\Gamma(Q)=\min_{\pi}{\mathcal{H}}(Q,\pi)$. We then look at the regularization term, which causes all the aforementioned challenges, and write:
@@
$=(Q(s,a)-\overline{r}(s,a))^{2}+(\mathbb{E}_{s^{\prime}}[V^{\pi}(s^{\prime})])^{2}+$
$\hskip 14.45377pt2(\overline{r}(s,a)-Q(s,a))\mathbb{E}_{s^{\prime}}[V^{\pi}(s^{\prime})]$
@@
^$\widehat{{\mathcal{H}}}(Q,\pi)\stackrel{{\scriptstyle def}}{{=}}\sum_{i\in[N]}w_{i}\mathbb{E}_{\rho^{i}}[{\mathcal{T}}^{\pi}[Q](s,a))]$
$-(\mathbb{E}_{\rho_{\pi}}[{\mathcal{T}}^{\pi}[Q](s,a))]-{\mathbb{E}_{\rho_{\pi}}[\log\pi(s,a)]})$
$-\alpha\mathbb{E}_{\rho^{u}}\Big{[}(Q(s,a)-\overline{r}(s,a))^{2}+(\mathbb{E}_{s^{\prime}}V^{\pi}(s^{\prime}))^{2}$
$+2\text{ReLU}(\overline{r}(s,a)-Q(s,a))\mathbb{E}_{s^{\prime}}V^{\pi}(s^{\prime})\Big{]}$ (7)
Let $\widehat{\Gamma}(Q)=\min_{\pi}\widehat{{\mathcal{H}}}(Q,\pi))$. The following proposition indicates that $\widehat{\Gamma}(Q)$ always lower-bounds $\Gamma(Q)$.
@@
We note that assuming $Q\geq 0$ is not restrictive, as if the expert’s rewards $r^{*}(s,a)$ are non-negative (typically the case), then the soft-Q function should also be non-negative. This is due to the fact that $Q(s,a)=\mathbb{E}[\sum_{{s_{t},a_{t}}}\gamma^{t}(r(s,a)-\log\pi(s,a))|(s_{0},a_{0})=(s,a)]\geq 0$. As $\widehat{\Gamma}(Q)$ provides a lower-bound approximation of $\Gamma(Q)$, maximizing $\widehat{\Gamma}(Q)$ over the $Q$-space would drive $\Gamma(Q)$ towards its maximum value. It is important to note that, given that the inner problem involves minimization, obtaining an upper-bound approximation function is easier. However, since the outer problem is a maximization one, an upper bound would not be helpful in guiding the resulting solution towards optimal ones.
The following theorem indicates that, in fact, the approximation $\widehat{\Gamma}(Q)$ is more tractable to use.
@@
For any $Q\geq 0$, the following results hold
  • (i)
[…]
  • (ii)
  • (iii)
    $\widehat{\Gamma}(Q)$ is concave for $Q\geq 0$
@@
The above theorem tells us that new objective $\widehat{\Gamma}(Q)$ has a closed form where $V^{\pi}(s)$ are replaced by $V^{Q}(s)$. Moreover $\widehat{\Gamma}(Q)$ is concave for all $Q\geq 0$. The concavity is particularly advantageous, as it guarantees that the optimization objective is well-behaved and has a unique solution $Q^{*}$ such that $(Q^{*},\pi^{Q^{*}})$ form a unique saddle point of $\max_{Q}\min_{\pi}~{}~{}\widehat{{\mathcal{H}}}(Q,\pi)$.
4 Practical Algorithm – SubIQ
The term $\mathbb{E}_{\rho_{\pi}}[{\mathcal{T}}^{\pi}[Q](s,a))]-{\mathbb{E}_{\rho_{\pi}}[\log\pi(s,a)]}$ can be written as (Garg et al., 2021):
$\mathbb{E}_{\rho_{\pi}}[{\mathcal{T}}^{\pi}[Q](s,a))]\!-\!{\mathbb{E}_{\rho_{\pi}}[\log\pi(s,a)]}\!=\!(1\!-\!\gamma)\mathbb{E}_{s_{0}}[V^{\pi}(s_{0})]$
$=\mathbb{E}_{\rho^{*}}[V^{\pi}(s)-\gamma\mathbb{E}_{s^{\prime}}V^{\pi}(s^{\prime})]$
$\mathbb{E}_{\rho_{\pi}}[{\mathcal{T}}^{\pi}[Q](s,a))]-{\mathbb{E}_{\rho_{\pi}}[\log\pi(s,a)]}=$
$=\sum_{i\in[N]}w_{i}\mathbb{E}_{\rho^{i}}[V^{\pi}(s)-\gamma\mathbb{E}_{s^{\prime}}V^{\pi}(s^{\prime})]$ (8)
$\hskip 14.45377pt-\sum_{i\in[N]}\mathbb{E}_{\rho^{i}}[V^{Q}(s)-\gamma\mathbb{E}_{s^{\prime}}V^{Q}(s^{\prime})]$
$\hskip 14.45377pt-\alpha\sum_{i\in[N]}w_{i}\mathbb{E}_{\rho^{i}}[\Delta^{Q}_{\overline{r}}(s,a)]$ (9)
where $\Delta^{Q}_{\overline{r}}(s,a)=(Q(s,a)-\overline{r}(s,a))^{2}+(\mathbb{E}_{s^{\prime}}V^{Q}(s^{\prime}))^{2}+2\text{ReLU}(\overline{r}(s,a)-Q(s,a))\mathbb{E}_{s^{\prime}}V^{Q}(s).$
Empirical expectation estimates.
In an empirical offline implementation, to maximize ^$\widehat{{\mathcal{H}}}(Q)$, samples $(s,a,s^{\prime})$ from demonstrations can be used to approximate the expectations over $\rho^{i}$: $\sum_{(s,a,s^{\prime})\in{\mathcal{D}}^{i}}[Q(s,a)-\gamma V^{Q}(s^{\prime})]$, $\sum_{(s,a,s^{\prime})\in{\mathcal{D}}^{i}}[V^{Q}(s)-\gamma V^{Q}(s^{\prime})]$, and $\sum_{(s,a)\in{\mathcal{D}}^{i}}[\Delta^{Q}_{\overline{r}}(s,a)]$. We note that in continuous-action controls, the computation of $V^{Q}(s)$ involves a sum over infinitely many actions, which is impractical. In this case, we can update both $Q$ and $\pi$ in a soft actor-critic (SAC) manner. That is, for each $\pi$, we update $Q$ towards $\max_{Q}\{\widehat{{\mathcal{H}}}(Q,\pi)\}$ and for each $Q$, we update $\pi$ to bring it towards $\pi^{Q}$ by solving $\max_{\pi}\{V^{\pi}(s)\}$, for all $s$. As shown above, ^$\widehat{{\mathcal{H}}}(Q,\pi)$ in concave in $Q$ and $V^{\pi}(s)$ is convex in $\pi$, so we can expect this SAC will exhibit good behavior and stability.
Estimating the Reward Reference.
@@ 4 Practical Algorithm – SubIQ
The reward reference function $\overline{r}(s,a)$ is a critical component of our learning algorithm. In the following, we outline our approach to automatically infer such a reward function from demonstrations of different expertise levels. The general idea is to learn a function that assigns higher values to higher expert-level demonstrations. To achieve this, let us define $R(\tau)=\sum_{(s,a)\in\tau}\overline{r}(s,a)$ (i.e., accumulated reward of trajectory $\tau$). For two trajectories $\tau_{i},~{}\tau_{j}$, let $\tau_{i}\prec\tau_{j}$ denote that $\tau_{i}$ is lower in quality compared to $\tau_{j}$ (i.e., $\tau_{i}$ belongs to demonstrations from lower-expert policies, compared to $\tau_{j}$). We model the probability $P(\tau_{i}\prec\tau_{j})$ by a softmax-normalized (or binary logit) distribution, which gives:
$P(\tau_{i}\prec\tau_{j})=\frac{\exp(R(\tau_{j}))}{\exp(R(\tau_{i}))+\exp(R(\tau_{j}))}$
Intuitively, increasing $R(\tau_{i})$ will lower $P(\tau_{i}\prec\tau_{j})$ and $P(\tau_{i}\prec\tau_{j})<0.5$ if $R(\tau_{i})<R(\tau_{j})$. Having this, we use the following loss function:
$\min_{\overline{r}}\Bigg{\{}{\mathcal{L}}(\overline{r})$ $=\sum_{i\in[N]}\sum_{(s,a),(s^{\prime},a^{\prime})\in{\mathcal{D}}^{i}}(\overline{r}(s,a)-\overline{r}(s^{\prime},a^{\prime}))^{2}$
$-\sum_{\begin{subarray}{c}h,k\in[N],h<k\\ \tau_{i}\in{\mathcal{D}}^{h},\tau_{j}\in{\mathcal{D}}^{k}\end{subarray}}\ln P(\tau_{i}\prec\tau_{j})+\phi(\overline{r})\Bigg{\}}$ (12)
@@
Here, the first term of ${\mathcal{L}}(\overline{r})$ serves to guarantee that the reward reference values for $(s,a)$ pairs within the same demonstration group are similar, and the second term aims to increase the likelihood that the accumulated rewards of trajectories adhere to the expert-level order, and the last term is a convex regularizer. Importantly, it can be shown below that ${\mathcal{L}}(\overline{r})$ is convex in $\overline{r}$, making the learning well-behaved.
@@
In practice, one can model $\overline{r}(s,a)$ by a neural network of parameters $\theta$ and optimize ${\mathcal{L}}(\theta)$ over $\theta$-space.
Conservative soft-Q learning
Over-estimation is a common issue in offline Q-learning due to out-of-distribution actions and function approximation errors (Fujimoto et al., 2018). We also observe this in our IL context. To overcome this issue and enhance stability, we leverage the approach in (Kumar et al., 2020) to enhance our inverse soft-Q learning. The aim is to learn a conservative soft-Q function that lower-bounds its true value. We formulate the conservative inverse soft-Q objective as:
$\small\max_{Q}\left\{\widehat{{\mathcal{H}}}^{C}(Q,\pi)=-\beta\sum_{\begin{subarray}{c}s\sim{\mathcal{D}}\\ a\sim\mu(a|s)\end{subarray}}[Q(s,a)]+\widehat{{\mathcal{H}}}(Q,\pi)\right\}$ (13)
where $\mu(a|s)$ is a particular state-action distribution. We note that in (13), the conservative term is added to the objective function, while in the conservative Q-learning algorithm (Kumar et al., 2020), this term is added to the Bellman error objective of each Q-function update. This difference makes the theory developed in (Kumar et al., 2020) not applicable. In Proposition 4.2 below, we show that solving $\max_{Q}{\widehat{{\mathcal{H}}}^{C}(Q)}$ will always yield a Q-function that is a lower bound to the Q function obtained by solving $\max_{Q}{\widehat{{\mathcal{H}}}(Q,\pi)}$.

### v2

@@ 3.2 Regularization with Reference Reward
where $\alpha>0$ is a weight parameter for the reward regularizer term. With (4), the goal is to find a policy with an occupancy distribution that matches with the occupancy distribution of different expertise levels appropriately (characterized by the weights, $w_{i}$). Simultaneously, it ensures that the learning rewards are close to the pre-assigned rewards, aiming to guide the learning policy towards replicating expert demonstrations, while also learning from sub-optimal demonstrations.
@@
3.3 Concave Lower-bound on Inverse Soft-Q with Reward Regularizer
Even though (4) can be directly solved to recover rewards, prior research suggests that transforming (4) into the Q-space will enhance efficiency. We delve into this transformation approach in this section. As discussed in Section 2, there is a one-to-one mapping between any reward function $r$ and a function $Q$ in the Q-space. Thus, the maximin problem in (4) can be equivalently transformed as:
$\max_{Q}\min_{\pi}\Big{\{}{\mathcal{H}}(Q,\pi)\stackrel{{\scriptstyle def}}{{=}}\mathbb{E}_{\rho^{U}}[{\mathcal{T}}^{\pi}[Q](s,a))]$ $-\mathbb{E}_{\rho_{\pi}}[{\mathcal{T}}^{\pi}[Q](s,a))]+{\mathbb{E}_{\rho_{\pi}}[\log\pi(s,a)]}$
@@ 3.3 Concave Lower-bound on Inverse Soft-Q with Reward Regularizer
${\mathcal{T}}^{\pi}[Q](s,a)=Q(s,a)-\gamma\mathbb{E}_{s^{\prime}}[V^{\pi}(s^{\prime})],~{}V^{\pi}(s)=\mathbb{E}_{a\sim\pi(a|s)}[Q(s,a)-\log\pi(a|s)]$
In the context of single-expert-level, [13] demonstrated that the objective function in the $Q$-space as given in Equation 2 is concave in $Q$ and convex in $\pi$, implying that the maximin problem always has a unique saddle point solution. Unfortunately, this property does not hold in our case.
@@
${\mathcal{H}}(Q,\pi)$ (as defined in Equation 5) is concave in $Q$ but is not convex in $\pi$.
In general, we can see that the first and second term of (6) are convex in $Q$, but the reward regularizer term, which can be written as
$\alpha\mathbb{E}_{\rho^{U}}\Big{[}(Q(s,a)-\overline{r}(s,a)-\mathbb{E}_{s^{\prime}\sim P(.|s,a)}\mathbb{E}_{a^{\prime}\sim\pi(.|s^{\prime})}(Q(s^{\prime},a^{\prime})-\log\pi(s^{\prime},a^{\prime})))^{2}\Big{]},$
is not concave in $\pi$ (details are shown in the Appendix). The property indicated in Proposition 3.1 implies that the maximin problem within the $Q$-space $\max_{Q}\min_{\pi}J(Q,\pi)$ may not have a unique saddle point solution and would be more challenging to solve, compared to the original inverse IQ-learn problem.
Another key property of Equation 2 is with regards to the inner minimization problem over $\pi$, which yields a unique closed-form solution, enabling the transformation of the max-min problem into a non-adversarial concave maximization problem within the Q-space. The closed-form solution was given by $\pi^{Q}=\text{argmax}_{\pi}~{}~{}V^{\pi}(s)$ for all $s\in S$. Unfortunately, this result also does not hold with the new objective function in (6), as formally stated below:
@@
${\mathcal{H}}(Q,\pi)$ may not necessarily be minimized at $\pi^{*}$ such that $\pi^{*}=\text{argmax}_{\pi}~{}~{}V^{\pi}(s)$, for all $s\in S$.
To overcome the above challenges, our approach involves constructing a more tractable objective function that is a lower bound on the objective of (6). Let us first define $\Gamma(Q)=\min_{\pi}{\mathcal{H}}(Q,\pi)$. We then look at the regularization term, which causes all the aforementioned challenges, and write:
@@
$=(Q(s,a)-\overline{r}(s,a))^{2}+(\mathbb{E}_{s^{\prime}}[V^{\pi}(s^{\prime})])^{2}+2(\overline{r}(s,a)-Q(s,a))\mathbb{E}_{s^{\prime}}[V^{\pi}(s^{\prime})]$
@@
^$\widehat{{\mathcal{H}}}(Q,\pi)\stackrel{{\scriptstyle def}}{{=}}\sum_{i\in[N]}w_{i}\mathbb{E}_{\rho^{i}}[{\mathcal{T}}^{\pi}[Q](s,a))]-(\mathbb{E}_{\rho_{\pi}}[{\mathcal{T}}^{\pi}[Q](s,a))]-{\mathbb{E}_{\rho_{\pi}}[\log\pi(s,a)]})$
$-\alpha\mathbb{E}_{\rho^{U}}\Big{[}(Q(s,a)-\overline{r}(s,a))^{2}+(\mathbb{E}_{s^{\prime}}V^{\pi}(s^{\prime}))^{2}+2\text{ReLU}(\overline{r}(s,a)-Q(s,a))\mathbb{E}_{s^{\prime}}V^{\pi}(s^{\prime})\Big{]}$ (6)
Let $\widehat{\Gamma}(Q)=\min_{\pi}\widehat{{\mathcal{H}}}(Q,\pi))$. The proposition below shows that $\widehat{\Gamma}(Q)$ always lower-bounds $\Gamma(Q)$.
@@
We note that assuming $Q\geq 0$ is not restrictive, as if the expert’s rewards $r^{*}(s,a)$ are non-negative (typically the case), then the true soft-Q function, defined as $Q^{*}(s,a)=\mathbb{E}[\sum_{{s_{t},a_{t}}}\gamma^{t}(r^{*}(s,a)-\log\pi(s,a))|(s_{0},a_{0})=(s,a)]$, should also be non-negative, for any $\pi$. As $\widehat{\Gamma}(Q)$ provides a lower-bound approximation of $\Gamma(Q)$, maximizing $\widehat{\Gamma}(Q)$ over the $Q$-space would drive $\Gamma(Q)$ towards its maximum value. It is important to note that, given that the inner problem involves minimization, obtaining an upper-bound approximation function is easier. However, since the outer problem is a maximization one, an upper bound would not be helpful in guiding the resulting solution towards optimal ones. The following theorem indicates that $\widehat{\Gamma}(Q)$ is more tractable to use.
@@
For any $Q\geq 0$, the following results hold: (i) The inner minimization problem $\min_{\pi}\widehat{{\mathcal{H}}}(Q,\pi)$ has a unique optimal solution $\pi^{Q}$ such that $\pi^{Q}=\text{argmin}_{\pi}V^{\pi}(s)$ for all $s\in S$ and $\pi^{Q}(a|s)=\frac{\exp(Q(s,a))}{\sum_{a}\exp(Q(s,a))}$, (ii) $\max_{\pi}V^{\pi}(s)=\log(\sum_{a}\exp(Q(s,a)))\stackrel{{\scriptstyle def}}{{=}}V^{Q}(s)$, and (iii) $\widehat{\Gamma}(Q)$ is concave for $Q\geq 0$.
@@
The above theorem tells us that new objective $\widehat{\Gamma}(Q)$ has a closed form where $V^{\pi}(s)$ is replaced by $V^{Q}(s)$. Moreover $\widehat{\Gamma}(Q)$ is concave for all $Q\geq 0$. The concavity is particularly advantageous, as it guarantees that the optimization objective is well-behaved and has a unique solution $Q^{*}$ such that $(Q^{*},\pi^{Q^{*}})$ form a unique saddle point of $\max_{Q}\min_{\pi}~{}~{}\widehat{{\mathcal{H}}}(Q,\pi)$. Thus, our tractable objective has all the nice properties that the original IQ-Learn objective had, while being able to work for the offline case with multiple levels of expert trajectories and our reward regularizer.
Algorithm 1 SPRINQL: Inverse soft-Q Learning with Sub-optimal Demonstrations
0:$({\mathcal{D}}^{1},{\mathcal{D}}^{2},...,{\mathcal{D}}^{N}),(w_{1},w_{2},...,w_{N}),\overline{r}_{\eta},Q_{\psi},\pi_{\theta}$
1:# estimate reward reference function
2:for iteration $i...N_{e}$ do
3:    $d\leftarrow(d^{1},d^{1},...,d^{N})\sim({\mathcal{D}}^{1},{\mathcal{D}}^{2},...,{\mathcal{D}}^{N})$
4:    from dataset $d$, calculate ${\mathcal{L}}(\overline{r}_{\eta})$ by (7)
5:    $\eta\leftarrow\eta-\nabla_{\eta}{\mathcal{L}}(\overline{r}_{\eta})$
6:end for
7:# train SPRINQL
8:for iteration $i...N$  do
9:    $d\leftarrow(d^{1},d^{1},...,d^{N})\sim({\mathcal{D}}^{1},{\mathcal{D}}^{2},...,{\mathcal{D}}^{N})$
10:    # Update Q function
11:    from dataset $d$, calculate ^$\widehat{{\mathcal{H}}}^{C}(Q_{\psi},\pi_{\theta})$ by (8)
12:    $\psi\leftarrow\psi+\nabla_{\psi}\widehat{{\mathcal{H}}}^{C}(Q_{\psi},\pi_{\theta})$
13:    # Update policy for actor-critic
14:    $\theta\leftarrow\theta+\nabla\left[\mathbb{E}\begin{subarray}{c}s\sim d\\ a\sim\pi_{\theta}(a|s)\end{subarray}[Q_{\psi}(s,a)-\ln(\pi_{\theta}(a|s))]\right]$
15:end for
3.4 SPRINQL Algorithm
Algorithm 1 provides the overall SPRINQL algorithm. We first estimate the reference rewards in lines 2-6 of Algorithm 1 and the overall process is described in Section 3.4.1. Before we proceed to the overall training, we have to estimate the weights, $w_{i}$ (associated with the ranked demonstration sets) employed in ^$\hat{\cal H}(Q,\pi)$. We provide a description of this estimation procedure in Section 3.4.2. Finally, to enhance stability and mitigate over-estimation issues commonly encountered in offline Q-learning, we employ a conservative version of ^$\hat{\cal H}(Q,\pi)$ in lines 8-15 of the algorithm and is described in Section 3.4.3. Some other practical considerations are discussed in the appendix.
3.4.1 Estimating the Reference Reward
@@
We outline our approach to automatically infer the reference rewards $\overline{r}(s,a)$ from the ranked demonstrations. The general idea is to learn a function that assigns higher values to higher expert-level demonstrations. To achieve this, let us define $R(\tau)=\sum_{(s,a)\in\tau}\overline{r}(s,a)$ (i.e., accumulated reward of trajectory $\tau$). For two trajectories $\tau_{i},~{}\tau_{j}$, let $\tau_{i}\prec\tau_{j}$ denote that $\tau_{i}$ is lower in quality compared to $\tau_{j}$ (i.e., $\tau_{i}$ belongs to demonstrations from lower-expert policies, compared to $\tau_{j}$). We follow the Bradley-Terry model of preferences [4, 5] to model the probability $P(\tau_{i}\prec\tau_{j})$ as $P(\tau_{i}\prec\tau_{j})=\frac{\exp(R(\tau_{j}))}{\exp(R(\tau_{i}))+\exp(R(\tau_{j}))}$ and use the following loss function:
$\min_{\overline{r}}\{{\mathcal{L}}(\overline{r})$ $=\sum_{i\in[N]}\sum_{(s,a),(s^{\prime},a^{\prime})\in{\mathcal{D}}^{i}}(\overline{r}(s,a)-\overline{r}(s^{\prime},a^{\prime}))^{2}-\!\!\!\!\sum_{\begin{subarray}{c}h,k\in[N],h>k,\tau_{i}\in{\mathcal{D}}^{h},\tau_{j}\in{\mathcal{D}}^{k}\end{subarray}}\!\!\!\!\ln P(\tau_{i}\prec\tau_{j})\}$ (7)
@@
where the first term of ${\mathcal{L}}(\overline{r})$ serves to guarantee that the reward reference values for $(s,a)$ pairs within the same demonstration group are similar, and the second term aims to increase the likelihood that the accumulated rewards of trajectories adhere to the expert-level order. Importantly, it can be shown below that ${\mathcal{L}}(\overline{r})$ is convex in $\overline{r}$ (Proposition 3.5), making the learning well-behaved. In practice, one can model $\overline{r}(s,a)$ by a neural network of parameters $\theta$ and optimize ${\mathcal{L}}(\theta)$ over $\theta$-space.
@@
3.4.2 Preference-based Weight Learning for $w_{i}$
Each weight parameter $w_{i}$ used in (4) should reflect the quality of the corresponding demonstration set ${\mathcal{D}}^{i}$, which can be evaluated by estimating the average expert rewards of these sets. Although this information is not directly available in our setting, the reward reference values discussed earlier provide a convenient and natural way to estimate them. This leads to the following formulation for inferring the weights $w_{i}$ from the ranked data: $w_{i}=\frac{\mathbb{E}_{(s,a)\sim D^{i}}\left[\bar{r}(s,a)\right]}{\sum_{j\in[N]}\mathbb{E}_{(s,a)\sim D^{j}}\left[\bar{r}(s,a)\right]}.$
3.4.3 Conservative soft-Q learning
Over-estimation is a common issue in offline Q-learning due to out-of-distribution actions and function approximation errors [11]. We also observe this in our IL context. To overcome this issue and enhance stability, we leverage the approach in [23] to enhance our inverse soft-Q learning. The aim is to learn a conservative soft-Q function that lower-bounds its true value. We formulate the conservative inverse soft-Q objective as:
^$\widehat{{\mathcal{H}}}^{C}(Q,\pi)=-\beta\sum_{\begin{subarray}{c}s\sim{\mathcal{D}},~{}a\sim\mu(a|s)\end{subarray}}[Q(s,a)]+\widehat{{\mathcal{H}}}(Q,\pi)$ (8)
where $\mu(a|s)$ is a particular state-action distribution. We note that in (8), the conservative term is added to the objective function, while in the conservative Q-learning algorithm [23], this term is added to the Bellman error objective of each Q-function update. This difference makes the theory developed in [23] not applicable. In Proposition 3.6 below, we show that solving $\max_{Q}{\widehat{{\mathcal{H}}}^{C}(Q)}$ will always yield a Q-function that is a lower bound to the Q function obtained by solving $\max_{Q}{\widehat{{\mathcal{H}}}(Q,\pi)}$.